# Blockchain-Based Ad Auctions and Bayesian Persuasion: An Analysis of Advertiser Behavior

### v1

@@ 4 Model Development
4.1 Basic Setup
We consider a blockchain-based advertising platform where $N$ advertisers compete in an auction for a single ad slot to be displayed to a user. The platform has private information about the user engagement level, denoted by $\theta\in\Theta$, where $\Theta$ is a finite set of possible engagement states (e.g., low, medium, high). The prior probability distribution over $\Theta$ is given by $\pi(\theta)=\Pr(\theta)$.
4.2 Advertisers
Each advertiser $i\in\mathcal{A}=\{1,2,...,N\}$ aims to maximize their expected utility, which depends on their valuation of the ad impression, $v_{i}(\theta)$, which is a function of the user engagement level $\theta$ and the bidding strategy of other advertisers.
@@ 4.3 Auction Mechanism
The platform conducts a sealed-bid, second-price auction:
  • Bidding: Advertisers submit bids $b_{i}\in\mathbb{R}_{+}$ without knowledge of others’ bids.
  • Allocation: The highest bidder wins the ad slot.
  • Payment: The winner pays the amount of the second-highest bid.
In a blockchain context, bids and outcomes are recorded transparently, but we assume the user engagement level $\theta$ remains private information to the platform.
@@ 4.4 Platform’s Signaling Strategy
The platform commits to a signaling policy $\sigma:\Theta\rightarrow\Delta(S)$, where:
  • $S$ is the set of possible signals the platform can send.
  • $\Delta(S)$ is the set of probability distributions over $S$.
@@ 4.6 Advertisers’ Bidding Strategy
where:
  • $c_{i}(b_{i},b_{-i})$ is the expected payment if advertiser $i$ wins.
  • $q_{i}(b_{i},b_{-i})$ is the probability that advertiser $i$ wins the auction.
In a symmetric equilibrium, advertisers adopt bidding strategies that are functions of their expected valuations.
@@ 5.2 Framework for Machine Learning Integration
We integrate machine learning into our methodology through three key steps: data collection and preparation; model selection and training; and prediction and integration with Bayesian persuasion.
5.2.1 Data Sources
We use data from ad auctions, which includes advertiser bids $b_{i}$ for various auctions, signals $s$ sent by the platform in those auctions, user engagement levels $\theta$, and auction outcomes. Additionally, we collect advertiser-specific features such as budget, industry sector, and bidding aggressiveness, as well as contextual features like time of day and ad category.
5.2.2 Data Privacy and Compliance
Given the sensitive nature of the data, we must ensure compliance with data protection regulations such as GDPR and CCPA. Anonymization and aggregation techniques should be employed to protect individual advertiser identities, thereby safeguarding privacy while allowing for effective model training.
5.2.3 Data Preprocessing
Data preprocessing involves several steps. First, we perform data cleaning by removing or imputing missing or erroneous data. Next, we engage in feature engineering to create relevant features that capture important aspects of advertiser behavior. We then normalize the features to ensure consistent ranges across variables, facilitating effective learning by the model. Finally, we split the dataset into training, validation, and test sets to enable proper evaluation and prevent overfitting.
@@ 5.3.2 Model Training
5.3.3 Model Evaluation
We evaluate the model’s performance using metrics such as Mean Absolute Error (MAE), which measures the average magnitude of errors without considering their direction; Root Mean Squared Error (RMSE), which penalizes larger errors more than MAE by squaring the errors before averaging and taking the square root; and R-squared ($R^{2}$), which represents the proportion of variance in the dependent variable that is predictable from the independent variables. These metrics provide a comprehensive understanding of the model’s accuracy and reliability in predicting advertiser bids.
@@ 6.5.1 Full Disclosure
  • Signals: $S=\{s_{L},s_{H}\}$, where $s_{L}$ corresponds to $\theta_{L}$ and $s_{H}$ corresponds to $\theta_{H}$.
  • Beliefs: $\mu(\theta_{L}\mid s_{L})=1$, $\mu(\theta_{H}\mid s_{H})=1$.
  • Advertisers’ Bids: $b_{i}^{*}(s_{L})=v_{i}(\theta_{L})$, $b_{i}^{*}(s_{H})=v_{i}(\theta_{H})$.
@@ 6.5.2 No Disclosure
  • Signals: $S=\{s\}$; the same signal is sent regardless of $\theta$.
  • Beliefs: $\mu(\theta_{L}\mid s)=\pi_{L}$, $\mu(\theta_{H}\mid s)=\pi_{H}$.
  • Advertisers’ Bids: $b_{i}^{*}(s)=\pi_{L}v_{i}(\theta_{L})+\pi_{H}v_{i}(\theta_{H})$.
@@ 6.5.3 Partial Disclosure
The platform can randomize signals to create mixed beliefs. For example:
  • Signals: $S=\{s_{1},s_{2}\}$.
  • Signaling Policy:

### v2

@@ 4 Model Development
We consider a blockchain-based advertising platform where $N$ advertisers compete in an auction for a single ad slot to be displayed to a user. The platform has private information about the user engagement level, denoted by $\theta\in\Theta$, where $\Theta$ is a finite set of possible engagement states (e.g., low, medium, high). The prior probability distribution over $\Theta$ is given by $\pi(\theta)=\Pr(\theta)$. Each advertiser $i\in\mathcal{A}=\{1,2,...,N\}$ aims to maximize their expected utility, which depends on their valuation of the ad impression, $v_{i}(\theta)$, which is a function of the user engagement level $\theta$ and the bidding strategy of other advertisers.
@@ 4.1 Auction Mechanism
The auction mechanism is a sealed-bid, second-price auction where advertisers submit bids without knowing others’ bids. The highest bidder wins the slot and pays the amount of the second-highest bid. While bids and outcomes are transparently recorded on the blockchain, the engagement level $\theta$ remains private to the platform.
@@ 4.2 Platform’s Signaling Strategy
The platform commits to a signaling policy $\sigma:\Theta\rightarrow\Delta(S)$, where $S$ is the set of possible signals the platform can send. $\Delta(S)$ is the set of probability distributions over $S$.
@@ 4.4 Advertisers’ Bidding Strategy
where $c_{i}(b_{i},b_{-i})$ is the expected payment if advertiser $i$ wins. $q_{i}(b_{i},b_{-i})$ is the probability that advertiser $i$ wins the auction. In a symmetric equilibrium, advertisers adopt bidding strategies that are functions of their expected valuations.
@@ 5.1 Framework for Machine Learning Integration
We integrate machine learning into our methodology through three key steps: data collection and preparation; model selection and training; and prediction and integration with Bayesian persuasion. We use data from ad auctions, which includes advertiser bids $b_{i}$ for various auctions, signals $s$ sent by the platform in those auctions, user engagement levels $\theta$, and auction outcomes. Additionally, we collect advertiser-specific features such as budget, industry sector, and bidding aggressiveness, as well as contextual features like time of day and ad category. Given the sensitive nature of the data, we must ensure compliance with data protection regulations such as GDPR and CCPA. Anonymization and aggregation techniques should be employed to protect individual advertiser identities, thereby safeguarding privacy while allowing for effective model training. Data preprocessing involves several steps. First, we perform data cleaning by removing or imputing missing or erroneous data. Next, we engage in feature engineering to create relevant features that capture important aspects of advertiser behavior. We then normalize the features to ensure consistent ranges across variables, facilitating effective learning by the model. Finally, we split the dataset into training, validation, and test sets to enable proper evaluation and prevent overfitting.
@@ 5.2 Model Selection and Training
We then evaluate the model’s performance using metrics such as Mean Absolute Error (MAE), which measures the average magnitude of errors without considering their direction; Root Mean Squared Error (RMSE), which penalizes larger errors more than MAE by squaring the errors before averaging and taking the square root; and R-squared ($R^{2}$), which represents the proportion of variance in the dependent variable that is predictable from the independent variables. These metrics provide a comprehensive understanding of the model’s accuracy and reliability in predicting advertiser bids.
@@ 6.5.1 Full Disclosure
Signals: $S=\{s_{L},s_{H}\}$, where $s_{L}$ corresponds to $\theta_{L}$ and $s_{H}$ corresponds to $\theta_{H}$. Beliefs: $\mu(\theta_{L}\mid s_{L})=1$, $\mu(\theta_{H}\mid s_{H})=1$. Advertisers’ Bids: $b_{i}^{*}(s_{L})=v_{i}(\theta_{L})$, $b_{i}^{*}(s_{H})=v_{i}(\theta_{H})$.
@@ 6.5.2 No Disclosure
Signals: $S=\{s\}$; the same signal is sent regardless of $\theta$. Beliefs: $\mu(\theta_{L}\mid s)=\pi_{L}$, $\mu(\theta_{H}\mid s)=\pi_{H}$. Advertisers’ Bids: $b_{i}^{*}(s)=\pi_{L}v_{i}(\theta_{L})+\pi_{H}v_{i}(\theta_{H})$.
@@ 6.5.3 Partial Disclosure
The platform can randomize signals to create mixed beliefs. Signals: $S=\{s_{1},s_{2}\}$. Signaling Policy: